### 28 Feb 2026 v15

# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

Abstract.
Training logistic regression over encrypted data has emerged as a prominent approach to addressing security concerns in recent years. In this paper, we introduce an efficient gradient variant, termed the quadratic gradient, which is specifically designed for privacy-preserving logistic regression while remaining equally effective in plaintext optimization. By incorporating this quadratic gradient, we enhance Nesterov’s Accelerated Gradient (NAG), Adaptive Gradient (AdaGrad), and Adam algorithms. We evaluate these enhanced algorithms across various datasets, with experimental results demonstrating state-of-the-art convergence rates that significantly outperform traditional first-order gradient methods. Furthermore, we apply the enhanced NAG method to implement homomorphic logistic regression training, achieving comparable performance within only four iterations. The proposed quadratic-gradient approach offers a unified framework that synergizes the advantages of first-order gradient methods and second-order Newton-type methods, suggesting broad applicability to diverse numerical optimization tasks.
Homomorphic Encryption, Logistic Regression, Quadratic Gradient, Simplified Fixed Hessian, Nesterov’s Accelerated Gradient
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
1.1. Background
Given a patient’s longitudinal healthcare data, one can train a logistic regression (LR) model to predict the likelihood of disease onset. However, personal health information (PHI) is inherently sensitive, and privacy concerns remain a significant barrier to the large-scale sharing and aggregation of biomedical data. To address this, Homomorphic Encryption (HE) offers a robust security paradigm: data can be transformed into ciphertexts and outsourced to the cloud for computation without ever being decrypted, thereby preventing unauthorized access by the service provider.
The annual iDASH competition has become a benchmark for evaluating such cryptographic schemes in biological contexts. Since 2014, iDASH has focused on genomic and biomedical privacy; specifically, the 2017 (Track 3) and 2018 (Track 2) competitions were dedicated to the development of high-performance HE-based solutions for training LR models over encrypted datasets.
1.2. Related work
Several studies have explored the implementation of logistic regression (LR) models within homomorphic encryption (HE) frameworks. Kim et al. (kim2018secure) addressed the challenges of encrypted LR training by employing full-batch gradient descent and utilizing least-squares approximation for the sigmoid function. During the 2017 iDASH competition, this problem was further investigated by Bonte and Vercauteren (IDASH2018bonte), Kim et al. (IDASH2018Andrey), Chen et al. (IDASH2018chen), and Crawford et al. (IDASH2018gentry). In the subsequent 2018 iDASH competition, Kim et al. (IDASH2019kim) and Blatt et al. (IDASH2019blatt) introduced efficient ciphertext packing techniques and semi-parallel algorithms to enhance computational throughput.
While other related works (kim2019secure; ogilvie2020improved) have examined various facets of encrypted optimization, the studies most pertinent to our research are (IDASH2018bonte) and (IDASH2018Andrey). Specifically, Bonte and Vercauteren (IDASH2018bonte) proposed a practical algorithm known as the Simplified Fixed Hessian (SFH) method. Our work extends the SFH approach and integrates the ciphertext packing mechanism proposed by Kim et al. (IDASH2018Andrey) to achieve high-efficiency homomorphic computation.
1.3. Contributions
Our specific contributions in this paper are summarized as follows:
  1. (1)
    We introduce a novel gradient variant, termed the quadratic gradient, which bridges first-order gradient-based methods and the second-order Newton-type algorithms. This enables a unified framework that synergistically leverages the computational efficiency of first-order methods and the rapid convergence of second-order approaches.
  2. (2)
    We develop three enhanced optimization algorithms by integrating the quadratic gradient into existing frameworks. Experimental results demonstrate that these enhanced algorithms achieve state-of-the-art performance in terms of convergence rates across diverse datasets.
  3. (3)
    We implement privacy-preserving logistic regression training using the enhanced Nesterov’s Accelerated Gradient (NAG) method. To the best of our knowledge, our implementation strikes an optimal balance between computational efficiency and storage overhead without significant performance degradation.
  4. (4)
    We propose a systematic framework that formalizes the derivation of constant Hessian approximations in fixed-Hessian methods (bohning1988monotonicity; IDASH2018bonte). Furthermore, we evaluate the practicality of employing such fixed-Hessian matrices to enhance computational efficiency in privacy-preserving contexts.
2. Preliminaries
In the following discussion, we employ square brackets $\cdot$] to denote the indexing of vectors and matrices. Specifically, for a vector $\boldsymbol{v}\in\mathbb{R}^{n}$ and a matrix $\mathbf{M}\in\mathbb{R}^{m\times n}$, $\boldsymbol{v}[i]$ (or $\boldsymbol{v}_{i}$) denotes the $i$-th element of $\boldsymbol{v}$, while $\mathbf{M}[i][j]$ (or $\mathbf{M}_{i,j}$) refers to the element in the $i$-th row and $j$-th column of $\mathbf{M}$.
2.1. Fully Homomorphic Encryption
Fully Homomorphic Encryption (FHE) is a cryptographic primitive that enables an arbitrary number of additions and multiplications directly on ciphertexts. While Gentry proposed the first feasible FHE scheme using a bootstrapping operation in 2009 (gentry2009fully), such schemes remain computationally intensive. Beyond the inherent complexity of FHE, the efficiency of homomorphic computation is significantly influenced by the choice of dataset encoding and the management of plaintext magnitudes (jaschke2016accelerating). To address the latter, Cheon et al. (cheon2017homomorphic) introduced an HE scheme featuring a rescaling procedure that effectively mitigates this technical bottleneck.
In our implementation of homomorphic LR algorithms, we adopt the open-source library HEAAN, which realizes this rescaling functionality. Furthermore, leveraging ciphertext packing is essential to achieve a superior amortized runtime. HEAAN supports Single Instruction, Multiple Data (SIMD) operations (smart2014fully), allowing multiple complex numbers to be packed into the slots of a single ciphertext polynomial and enabling efficient rotation operations across these slots. The mathematical foundations of the HEAAN scheme are detailed in (IDASH2018Andrey; kim2018secure; han2018efficient), with the underlying abstract algebra further elaborated in (artin2011algebra).
2.2. Database Encoding Method
Kim et al. (IDASH2018Andrey) proposed a highly efficient database encoding method that leverages SIMD batching to optimize both computational and storage resources. Given a training dataset $Z$ containing $n$ samples, each with $d+1$ covariates, the authors packed the entire matrix into a single ciphertext in a row-major format, thereby maximizing the utilization of available ciphertext slots.
Under this encoding paradigm, the data matrix $Z$ can be manipulated directly through homomorphic operations on its ciphertext representation, $\text{Enc}(Z)$. This is achieved using a minimal set of primitive operations: rotation, addition, and multiplication. For example, to isolate the first column of $\text{Enc}(Z)$ while masking others, one can multiply the ciphertext by a constant matrix $F$ (where the first column contains ones and all other elements are zero).
Han et al. (han2018efficient) delineated several fundamental yet critical operations employed in the implementation by Kim et al. (IDASH2018Andrey), such as the SumColVec procedure for computing column-wise summations of a matrix. By composing these primitive operations, more sophisticated computations—most notably the gradient evaluation required for logistic regression—can be efficiently realized in the encrypted domain.
2.3. Logistic Regression Model
Logistic regression (LR) is widely used in binary classification tasks to infer whether a binary-valued variable belongs to a certain class or not. LR generalizes linear regression (murphy2012machine) by mapping the linear predictor $\boldsymbol{\beta}^{\top}\mathbf{x}$ to the interval $(0,1)$ via the sigmoid function $\sigma(z)=(1+\exp(-z))^{-1}$. Here, $\boldsymbol{\beta}\in\mathbb{R}^{d+1}$ denotes the model parameters, and $\mathbf{x}=(1,x_{1},\ldots,x_{d})^{\top}\in\mathbb{R}^{d+1}$ represents the input covariate vector. For a class label $y\in\{+1,-1\}$, the conditional probabilities are formulated as:
$\Pr(y=+1\mid\mathbf{x},\boldsymbol{\beta})$ $=\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})=\frac{1}{1+e^{-\boldsymbol{\beta}^{\top}\mathbf{x}}},$
$\Pr(y=-1\mid\mathbf{x},\boldsymbol{\beta})$ $=1-\sigma(\boldsymbol{\beta}^{\top}\mathbf{x})=\frac{1}{1+e^{\boldsymbol{\beta}^{\top}\mathbf{x}}}.$
A classification decision is typically made by comparing the output probability against a predefined threshold (e.g., $0.5$).
Training an LR model is framed as a Maximum Likelihood Estimation (MLE) problem, seeking $\boldsymbol{\beta}$ that maximizes the likelihood $L(\boldsymbol{\beta})=\prod_{i=1}^{n}\Pr(y_{i}\mid\mathbf{x}_{i},\boldsymbol{\beta})$. For computational tractability, we maximize the log-likelihood function $l(\boldsymbol{\beta})$:
$l(\boldsymbol{\beta})=\ln L(\boldsymbol{\beta})=-\sum_{i=1}^{n}\ln(1+e^{-y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}}),$
where $n$ denotes the number of training samples. Since $l(\boldsymbol{\beta})$ lacks a closed-form solution, parameters are typically estimated using iterative optimization: (a) first-order methods, such as gradient descent; and (b) second-order methods, such as Newton’s method. The gradient and Hessian of $l(\boldsymbol{\beta})$ are given by:
$\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ $=\sum_{i=1}^{n}(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))y_{i}\mathbf{x}_{i},$
$\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$ $=\sum_{i=1}^{n}(y_{i}\mathbf{x}_{i})\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})-1)(y_{i}\mathbf{x}_{i})^{\top}=-\mathbf{X}^{\top}\mathbf{S}\mathbf{X},$
where $\mathbf{S}$ is a diagonal matrix with entries $\mathbf{S}_{ii}=\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i})(1-\sigma(y_{i}\boldsymbol{\beta}^{\top}\mathbf{x}_{i}))$, and $\mathbf{X}\in\mathbb{R}^{n\times(d+1)}$ denotes the dataset matrix.
The log-likelihood $l(\boldsymbol{\beta})$ is a concave function with at most one unique global maximum (Allison2008LRConvergenceFail), where its gradient is zero. Newton’s method, as a second-order technique, finds the optimal $\boldsymbol{\beta}$ by iteratively solving for the roots of the optimality condition $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})=0$.
3. Technical Details
Newton’s method requires recomputing the Hessian matrix and its inverse at each iteration, a process that is often computationally prohibitive. To mitigate this burden, the fixed Hessian Newton’s method (also known as the lower bound principle) replaces the varying Hessian with a constant matrix $\bar{H}$, thereby enhancing computational efficiency.
Böhning and Lindsay (bohning1988monotonicity) established that the convergence of Newton’s method is guaranteed provided that $\bar{H}\preceq\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})$, where $\bar{H}$ is a symmetric negative-definite matrix independent of $\boldsymbol{\beta}$. Here, $\preceq$” denotes the Loewner partial order, implying that the difference $\nabla_{\boldsymbol{\beta}}^{2}l(\boldsymbol{\beta})-\bar{H}$ is positive semi-definite. By employing such a fixed Hessian $\bar{H}$, the Newton iteration simplifies to the following update rule:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\bar{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}_{t}).$
Böhning and Lindsay further suggest that the fixed matrix $\bar{H}=-\frac{1}{4}\mathbf{X}^{\top}\mathbf{X}$ serves as an effective lower bound for the Hessian of the log-likelihood function $l(\boldsymbol{\beta})$ in logistic regression.
3.1. Simplified Fixed Hessian
Bonte and Vercauteren (IDASH2018bonte) further simplify the fixed Hessian $\bar{H}$ to facilitate its inversion within the encrypted domain. Specifically, they approximate $\bar{H}$ with a diagonal matrix $\mathbf{B}$, where each diagonal element $b_{ii}$ is defined as the sum of the corresponding row in $\bar{H}$. Furthermore, they propose an optimized computational schedule to construct $\mathbf{B}$ more efficiently. This diagonal approximation is formulated as follows:
$\mathbf{B}=\begin{bmatrix}\sum_{j=0}^{d}\bar{h}_{0j}&0&\ldots&0\\ 0&\sum_{j=0}^{d}\bar{h}_{1j}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&\sum_{j=0}^{d}\bar{h}_{dj}\end{bmatrix}$,
where $\bar{h}_{kj}$ denotes the $(k,j)$-th entry of $\bar{H}$.
This diagonal structure significantly simplifies the matrix representation, allowing $\mathbf{B}$ to be efficiently derived from $\bar{H}$. The inverse $\mathbf{B}^{-1}$ can be homomorphically approximated by computing the reciprocal of each diagonal element via an iterative Newton-Raphson method with a carefully chosen initial value. Consequently, the Simplified Fixed Hessian (SFH) update rule is formulated as:
$\boldsymbol{\beta}_{t+1}$ $=\boldsymbol{\beta}_{t}-\mathbf{B}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}_{t})$
$=\boldsymbol{\beta}_{t}-\begin{bmatrix}b_{00}^{-1}&0&\ldots&0\\ 0&b_{11}^{-1}&\ldots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\ldots&b_{dd}^{-1}\end{bmatrix}\begin{bmatrix}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\end{bmatrix}$
$=\boldsymbol{\beta}_{t}-\begin{bmatrix}b_{00}^{-1}\cdot\nabla_{0}\\ b_{11}^{-1}\cdot\nabla_{1}\\ \vdots\\ b_{dd}^{-1}\cdot\nabla_{d}\end{bmatrix}$,
where $b_{ii}^{-1}$ denotes the reciprocal of $\sum_{j=0}^{d}\bar{h}_{ij}$, and $\nabla_{i}$ represents the $i$-th element of the gradient vector $\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}_{t})$.
Consider a specific scenario where all diagonal elements $b_{ii}^{-1}$ share a constant value $\eta$ (with $\eta>0$). In this case, the iterative update for the SFH method reduces to:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-(-\eta)\begin{bmatrix}\nabla_{0}\\ \nabla_{1}\\ \vdots\\ \nabla_{d}\end{bmatrix}=\boldsymbol{\beta}_{t}+\eta\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}_{t}),$
which is mathematically equivalent to the vanilla gradient ascent update. This alignment provides the foundational insights for the unified framework proposed in this study, which reconciles fixed-Hessian methods with gradient-based optimization.
We characterize $\mathbf{B}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta})$ as an augmented gradient and incorporate a dedicated learning rate to capture its underlying dynamics. Provided this learning rate decays from a value exceeding unity (e.g., $2.0$) toward $1.0$ within a finite number of iterations, the fixed-Hessian principle theoretically ensures the convergence of our proposed scheme.
Despite its advantages, the SFH method (IDASH2018bonte) suffers from two primary limitations. First, its convergence is strictly guaranteed only when all entries of the symmetric matrix $\bar{H}$ are non-positive. While feature normalization (e.g., scaling to $[0,1]$) often satisfies this condition in machine learning, it may not hold in general numerical optimization. Second, both the simplified Hessian $\mathbf{B}$ and the fixed Hessian $\bar{H}=-\frac{1}{4}\mathbf{X}^{\top}\mathbf{X}$ are prone to singularity, particularly in high-dimensional sparse settings such as the MNIST dataset.
To address these issues, we generalize the SFH framework to ensure the invertibility of the simplified Hessian under arbitrary conditions. Building on this generalization, we introduce an accelerated gradient variant, termed the quadratic gradient.
3.2. Quadratic Gradient: Definition
Suppose a differentiable scalar-valued function $F(\mathbf{x})$ has a gradient $\boldsymbol{g}$ and a Hessian matrix $\mathbf{H}$. For the maximization task, let $\bar{\mathbf{H}}$ be any matrix satisfying the condition $\bar{\mathbf{H}}\preceq\mathbf{H}$ in the Loewner partial ordering. We represent these quantities as:
$\boldsymbol{g}=\begin{bmatrix}g_{0}\\ g_{1}\\ \vdots\\ g_{d}\end{bmatrix},\quad\mathbf{H}=\begin{bmatrix}\nabla_{00}^{2}&\nabla_{01}^{2}&\dots&\nabla_{0d}^{2}\\ \nabla_{10}^{2}&\nabla_{11}^{2}&\dots&\nabla_{1d}^{2}\\ \vdots&\vdots&\ddots&\vdots\\ \nabla_{d0}^{2}&\nabla_{d1}^{2}&\dots&\nabla_{dd}^{2}\end{bmatrix}$,
$\quad\quad\quad\bar{\mathbf{H}}=\begin{bmatrix}\bar{h}_{00}&\bar{h}_{01}&\dots&\bar{h}_{0d}\\ \bar{h}_{10}&\bar{h}_{11}&\dots&\bar{h}_{1d}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{h}_{d0}&\bar{h}_{d1}&\dots&\bar{h}_{dd}\end{bmatrix}$,
where $\nabla_{ij}^{2}=\frac{\partial^{2}F}{\partial x_{i}\partial x_{j}}$. We construct a diagonal matrix $\tilde{\mathbf{B}}$, referred to as the diagonal Hessian approximation, where each diagonal entry $\tilde{B}_{kk}$ is defined as:
$\tilde{B}_{kk}=-\epsilon-\sum_{i=0}^{d}|\bar{h}_{ki}|.$
Formally, $\tilde{\mathbf{B}}$ is structured as:
$\tilde{\mathbf{B}}=\begin{bmatrix}-\epsilon-\sum_{j=0}^{d}|\bar{h}_{0j}|&0&\dots&0\\ 0&-\epsilon-\sum_{j=0}^{d}|\bar{h}_{1j}|&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&-\epsilon-\sum_{j=0}^{d}|\bar{h}_{dj}|\end{bmatrix}$,
where $\epsilon>0$ is a small constant (typically set to $1\text{e}{-8}$) introduced for numerical stability, such as ensuring strict negative definiteness and preventing division by zero during inversion.
For the approximation $\tilde{B}$ to serve as a valid lower bound within the fixed-Hessian framework, it must satisfy the convergence criterion $\tilde{B}\leq H$. Given our foundational assumption that $\bar{H}\leq H$, this requirement is satisfied by establishing that $\tilde{B}\leq\bar{H}$. Following the analytical approach detailed in (IDASH2018bonte), it can be shown that $\tilde{B}$ consistently maintains this lower-bound property relative to $\bar{H}$, thereby ensuring stable convergence in the encrypted domain without the computational overhead of the full Hessian.
Lemma 3.1.
Let $\mathbf{A}\in\mathbb{R}^{n\times n}$ be a symmetric matrix, and let $\mathbf{B}$ be a diagonal matrix with entries $B_{kk}=-\epsilon-\sum_{i=1}^{n}|A_{ki}|$ for $k=1,\dots,n$. Then, $\mathbf{B}\preceq\mathbf{A}$.
Proof.
By the definition of Loewner ordering, it suffices to prove that the difference matrix $\mathbf{C}=\mathbf{A}-\mathbf{B}$ is positive definite ($\mathbf{C}\succ 0$). The entries of $\mathbf{C}$ are given by $C_{ii}=A_{ii}+\epsilon+\sum_{k=1}^{n}|A_{ik}|$ for $i=j$, and $C_{ij}=A_{ij}$ for $i\neq j$. According to Gerschgorin’s Circle Theorem, any eigenvalue $\lambda$ of $\mathbf{C}$ lies within at least one disk $|\lambda-C_{ii}|\leq\sum_{j\neq i}|C_{ij}|$. Thus,
$\lambda\geq C_{ii}-\sum_{j\neq i}|A_{ij}|=A_{ii}+\epsilon+|A_{ii}|+\sum_{j\neq i}|A_{ij}|-\sum_{j\neq i}|A_{ij}|=A_{ii}+|A_{ii}|+\epsilon.$
Since $A_{ii}+|A_{ii}|\geq 0$, we conclude that $\lambda\geq\epsilon>0$ for all eigenvalues $\lambda$. Thus, $\mathbf{C}$ is positive definite, implying $\mathbf{B}\preceq\mathbf{A}$. ∎
Definition 3.2 (Quadratic Gradient).
Utilizing the diagonal approximation $\tilde{\mathbf{B}}$, we define the quadratic gradient as $\mathbf{G}=\bar{\mathbf{B}}\boldsymbol{g}$, where $\bar{\mathbf{B}}$ is a diagonal matrix with entries $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$. For the maximization of $F(\mathbf{x})$, we employ the iterative update rule:
$\mathbf{x}_{t+1}=\mathbf{x}_{t}+N_{t}\cdot\mathbf{G},$
where $N_{t}\geq 1$ is a dynamic learning rate designed to decay toward $1.0$ over a finite number of iterations.
Minimizing $F(\mathbf{x})$ is equivalent to maximizing $F(\mathbf{x})$. In this context, $\tilde{\mathbf{B}}$ can be constructed using either a lower bound of the Hessian of $F(\mathbf{x})$ or an upper bound of the Hessian of $F(\mathbf{x})$. Notably, $\bar{\mathbf{H}}$ may be set to the Hessian $\mathbf{H}$ itself when available, bypassing the need for separate bound estimation.
3.2.1. Simplified Application-Oriented Version
While Newton’s method is fundamentally a root-finding algorithm for the optimality condition $\nabla F(\mathbf{x})=0$, its application depends on the specific optimization objective. In minimization scenarios, the Hessian $\mathbf{H}$ is typically positive definite, directing the update toward a local minimum. Conversely, for the maximization tasks addressed in this work, $\mathbf{H}$ is assumed to be negative definite.
When utilized in its standalone form, the quadratic gradient $\mathbf{G}$ functions as a fixed-Hessian Newton update. This provides a computationally efficient representation that unifies the convergence properties of second-order methods with the structural simplicity of gradient-based updates:
$\mathbf{x}_{t+1}=\mathbf{x}_{t}\pm N_{t}\cdot\mathbf{G},$
where the sign is set to ++ for maximization and - for minimization.
In practical applications where the Hessian matrix $\mathbf{H}$ is explicitly obtainable, we can bypass the formal construction of the quadratic gradient and derive it directly as:
$\mathbf{G}$ $=\text{diag}(\bar{B}_{00},\bar{B}_{11},\dots,\bar{B}_{dd})\cdot\boldsymbol{g}$
$=\begin{bmatrix}\bar{B}_{00}&0&\dots&0\\ 0&\bar{B}_{11}&\dots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\dots&\bar{B}_{dd}\end{bmatrix}\cdot\boldsymbol{g}$
(1a) $=\begin{bmatrix}\bar{B}_{00}\\ \bar{B}_{11}\\ \vdots\\ \bar{B}_{dd}\end{bmatrix}\odot\boldsymbol{g},$
where each diagonal element $\bar{B}_{ii}$ is defined as:
$\bar{B}_{ii}=\left(\epsilon+\sum_{j=0}^{d}|\nabla_{ij}^{2}|\right)^{-1}.$
Here, $\nabla_{ij}^{2}=\frac{\partial^{2}F(\mathbf{x})}{\partial x_{i}\partial x_{j}}$ Here, $\nabla_{ij}^{2}=\frac{\partial^{2}F(\mathbf{x})}{\partial x_{i}\partial x_{j}}$ denotes the $(i,j)$-th entry of the Hessian matrix itself, representing the second-order partial derivative of the objective function with respect to variables $x_{i}$ and $x_{j}$. The operator $\odot$” denotes the Hadamard product (i.e., element-wise multiplication). This reformulation transitions the computation from a standard matrix-vector product to a vectorized form that is highly compatible with HE-SIMD parallel processing. Specifically, by transforming the diagonal matrix $\text{diag}(\bar{B}_{00},\dots,\bar{B}_{dd})$ into a compact (plaintext) vector and computing its Hadamard product with the encrypted vector $\boldsymbol{g}$, the cloud server can execute the entire operation in a single homomorphic multiplication cycle. This approach optimizes both storage and execution time by circumventing the processing of off-diagonal zero elements.
In such cases, the surrogate Hessian matrix constructed still adheres to the convergence criteria of the fixed-Hessian Newton’s method. Consistent with our unified framework, the update rule employs an addition (++) sign for maximization tasks and a subtraction (-) sign for minimization tasks:
$\mathbf{x}_{t+1}=\mathbf{x}_{t}\pm N_{t}\cdot G,$
where the choice of sign ensures alignment with the respective optimization objective.
3.2.2. Systematic Fixed-Hessian Identification
Böhning and Lindsay (bohning1988monotonicity) did not provide a systematic methodology for constructing constant Hessian approximations, potentially due to the scarcity of such matrices for general objective functions. In this work, we introduce a structured procedure to identify and construct these fixed Hessian matrices. Our proposed approach serves as a diagnostic tool: identifying a fixed Hessian via this method provides sufficient evidence of its existence. However, it is important to note that our technique constitutes a sufficient but not necessary condition; thus, the inability to detect a fixed Hessian through our procedure does not preclude its existence.
Notably, our quadratic gradient approach does not strictly depend on a constant Hessian replacement. Instead, it allows for the use of the original, varying Hessian to construct the dynamic diagonal matrix $\tilde{\mathbf{B}}$. Under this more flexible framework, we present a methodical three-step approach to evaluate the feasibility of employing a fixed Hessian for computational efficiency:
  • Step 1: Construction. Derive the diagonal Hessian approximation $\tilde{\mathbf{B}}$ directly from the Hessian of the objective function.
  • Step 2: Optimization and Bound Identification. For each diagonal element $\tilde{B}_{kk}$, determine whether a constant maximum for $|\tilde{B}_{kk}|$ exists (equivalent to a constant minimum for $\bar{B}_{kk}=1/|\tilde{B}_{kk}|$). If each $|\tilde{B}_{kk}|$ achieves a global upper bound $|\tilde{M}_{kk}|$, then a constant diagonal matrix with entries $1/|\tilde{M}_{kk}|$ satisfies the convergence criteria for the SFH method.
  • Step 3: Performance Evaluation. Assess the practical performance of the resulting fixed Hessian. While such a replacement ensures theoretical convergence, it may not provide a tight bound to the true Hessian, potentially leading to a suboptimal convergence rate in practice.
This systematic search can be applied to re-derive the fixed Hessian matrix $\frac{1}{4}\mathbf{X}^{\top}\mathbf{X}$ originally proposed in (bohning1988monotonicity) for binary logistic regression, thereby unifying our framework with classical results.
Theoretically, the vanilla first-order gradient update, $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\alpha\boldsymbol{g}$ (where $\alpha$ is a sufficiently small step size), can be interpreted as a degenerate case of the fixed-Hessian Newton’s method. In this view, the inverse of the surrogate Hessian is simplified to a uniform scalar $\alpha\mathbf{I}$. However, such a scalar approximation fails to capture the non-uniform curvature of the objective function. Consequently, while it offers baseline stability, it lacks the adaptive convergence required for practical efficiency—particularly in high-dimensional encrypted optimization, where the high computational cost of each iteration necessitates a minimal number of steps to reach convergence.
3.2.3. Unitary Scalars as Degenerate Vectors
In numerical optimization, acceleration techniques for Newton-type methods primarily aim to enhance convergence robustness while mitigating computational overhead. A prevalent strategy is damping (or line search), where the full Newton step is modulated by a scaling factor to ensure global convergence. Mathematically, this damping parameter acts as a scalar multiplier, functionally analogous to the learning rate in first-order methods. However, unlike fixed step sizes, this parameter is adaptively determined to balance aggressive local progress with global stability, serving as a curvature-aware step controller.
From a structural perspective, the scalar parameter $\alpha$ in line-search methods can be interpreted as a degenerate vector where all entries are identical, i.e., $\alpha,\alpha,\dots,\alpha]^{\top}$. This representation reveals an inherent informational redundancy, where a scalar is essentially a high-dimensional vector with perfectly correlated components. In both first-order and second-order optimization, a standard learning rate or a line-search factor $\alpha$ imposes a uniform scaling across all dimensions of the update space. Under this interpretation, a scalar step size is not fundamentally distinct from a vector; rather, it represents a specific homogeneous structure where the update magnitude is constrained to be identical across all degrees of freedom.
We contend that the search space for Newton-type methods should ideally be represented by a search matrix $\mathbf{S}\in\mathbb{R}^{d\times d}$, which is isomorphic to the Hessian matrix $\mathbf{H}$. The general update rule is formulated as:
$\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}-\mathbf{S}\mathbf{H}^{-1}\nabla_{\boldsymbol{\beta}}l(\boldsymbol{\beta}_{t}).$
Standard line-search techniques drastically simplify this search space by substituting the full matrix $\mathbf{S}$ with a scalar $\alpha$. Consequently, this scalar acts as a degenerate identity matrix $\alpha\mathbf{I}$, collapsing the complexity of the directional update from $\mathcal{O}(d^{2})$ to $\mathcal{O}(1)$ and constraining the optimization trajectory to a single-degree-of-freedom manifold.
By expanding this collapsed scalar into a non-homogeneous vector, our proposed quadratic gradient framework restores the essential degrees of freedom required for efficient convergence in high-dimensional landscapes, agnostic of whether the optimization is conducted in the encrypted or plaintext domain. Consequently, we contend that Newton acceleration techniques relying on a single scalar, such as conventional line search, represent a degenerate special case of our quadratic gradient framework.
This evolutionary trajectory—moving from constrained scalar updates toward high-dimensional adaptive control—mirrors the historical development of first-order optimization: progressing from fixed step sizes (Vanilla SGD) to adaptive scalar rates (NAG), and ultimately to element-wise adaptive algorithms (e.g., AdaGrad, Adam). In the following sections, we demonstrate that the learning rate mechanism within our quadratic gradient framework undergoes a parallel transition. For instance, our Enhanced Adam algorithm can be viewed as a formal integration of the fixed-Hessian Newton method with a modernized, element-wise adaptive learning scheme.
Algorithm 1 The Enhanced Nesterov’s Accelerated Gradient Algorithm [^2] The Python implementation of Algorithm 2, optimized for Google Colab, is available at: https://anonymous.4open.science/r/IDASH2017-245B/Python3Experiments/.
1:Require: training dataset $X\in\mathbb{R}^{n\times(1+d)}$; training label $Y\in\mathbb{R}^{n\times 1}$; learning rate $lr\in\mathbb{R}$; and the number $\kappa$ of iterations;
2:Ensure: the parameter vector $V\in\mathbb{R}^{(1+d)}$
3:Set $\bar{H}\leftarrow-\frac{1}{4}X^{\top}X$ # $\bar{H}\in\mathbb{R}^{(1+d)\times(1+d)}$
4:Set $V\leftarrow\boldsymbol{0}$, $W\leftarrow\boldsymbol{0}$, $\bar{B}\leftarrow\boldsymbol{0}$ # $V\in\mathbb{R}^{(1+d)}$, $W\in\mathbb{R}^{(1+d)}$, $\bar{B}\in\mathbb{R}^{(1+d)\times(1+d)}$
5:for $i:=0$ to $d$ do
6:  # $\epsilon$ is a small positive constant such as 1e81e-8
7:  $\bar{B}[i][i]\leftarrow\epsilon$
8:  for $j:=0$ to $d$ do
9:   $\bar{B}[i][i]\leftarrow\bar{B}[i][i]+|\bar{H}[i][j]|$
10:  end for
11:end for
12:Set $alpha_{0}\leftarrow 0.01$
13:Set $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
14:for $count:=1$ to $\kappa$ do
15:  # $Z\in\mathbb{R}^{n}$ will store the inputs for Sigmoid function
16:  Set $Z\leftarrow\boldsymbol{0}$
17:  for $i:=1$ to $n$ do
18:   for $j:=0$ to $d$ do
19:    $Z[i]\leftarrow Z[i]+Y[i]\times V[j]\times X[i][j]$
20:   end for
21:  end for
22:  # $\boldsymbol{\sigma}\in\mathbb{R}^{n}$ will store the outputs of Sigmoid function
23:  Set $\boldsymbol{\sigma}\leftarrow\boldsymbol{0}$
24:  for $i:=1$ to $n$ do
25:   $\boldsymbol{\sigma}[i]\leftarrow 1/(1+\exp(-Z[i]))$
26:  end for
27:  Set $\boldsymbol{g}\leftarrow\boldsymbol{0}$
28:  for $j:=0$ to $d$ do
29:   for $i:=1$ to $n$ do
30:    $\boldsymbol{g}[j]\leftarrow\boldsymbol{g}[j]+(1-\boldsymbol{\sigma}[i])\times Y[i]\times X[i][j]$
31:   end for
32:  end for
33:  Set $G\leftarrow\boldsymbol{0}$
34:  for $j:=0$ to $d$ do
35:   $G[j]\leftarrow\bar{B}[j][j]\times\boldsymbol{g}[j]$
36:  end for
37:  Set $\eta\leftarrow(1-alpha_{0})/alpha_{1}$
38:  Set $\gamma\leftarrow lr/(n\times count)$
39:  for $j:=0$ to $d$ do
40:   $w_{temp}\leftarrow V[j]+(1+\gamma)\times G[j]$
41:   $V[j]\leftarrow(1-\eta)\times w_{temp}+\eta\times W[j]$
42:   $W[j]\leftarrow w_{temp}$
43:  end for
44:  $alpha_{0}\leftarrow alpha_{1}$
45:  $alpha_{1}\leftarrow 0.5\times(1+\sqrt{1+4\times alpha_{0}^{2}})$
46:end for
47:Return $V$
3.3. Quadratic Gradient: Algorithms
The quadratic gradient framework can be generalized to enhance various first-order optimization algorithms by incorporating second-order curvature information into their respective update rules:
  1. (1)
    Nesterov’s Accelerated Gradient (NAG): NAG is a variant of the momentum method designed to provide the momentum term with greater prescience. The standard iterative formulas for NAG in a gradient ascent context are:
    (2) $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\nabla J(\boldsymbol{\beta}_{t}),$
    (3) $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})V_{t+1}+\gamma_{t}V_{t},$
    where $V_{t+1}$ is an intermediate variable and $\gamma_{t}\in(0,1)$ is a smoothing parameter for the moving average, enabling gradient evaluation at an approximate future position (IDASH2018Andrey). The Enhanced NAG algorithm modifies the original update rule (2) by substituting the vanilla gradient with the quadratic gradient:
    $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\mathbf{B}^{-1}\nabla J(\boldsymbol{\beta}_{t})=\boldsymbol{\beta}_{t}+N_{t}\mathbf{G}_{t},$
    where $N_{t}$ is the adaptive learning rate for the enhanced variant. As detailed in Algorithm 2, we typically set $N_{t}=1+\eta_{t}$, where $\eta_{t}$ follows a specific decay schedule.
  2. (2)
    AdaGrad: AdaGrad is particularly well-suited for sparse data by adapting the learning rate for each parameter based on historical gradients. The update rules for standard AdaGrad and its Enhanced Quadratic version for each parameter $\boldsymbol{\beta}_{i}$ at iteration $t$ are given by:
    $\text{(Standard)}\quad\boldsymbol{\beta}_{i}^{(t+1)}$ $=\boldsymbol{\beta}_{i}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}(g_{i}^{(k)})^{2}}}\cdot g_{i}^{(t)},$
    $\text{(Enhanced)}\quad\boldsymbol{\beta}_{i}^{(t+1)}$ $=\boldsymbol{\beta}_{i}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}(G_{i}^{(k)})^{2}}}\cdot G_{i}^{(t)}.$
  3. (3)
    Adam: Adam integrates the benefits of AdaGrad’s element-wise adaptation with the stability of momentum, enabling faster convergence in the presence of sparse or noisy gradients. Following the same logic as the AdaGrad enhancement, the Enhanced Adam algorithm replaces the vanilla gradient with its quadratic form and employs the specialized learning rate within its bias-corrected first and second moment estimates.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 1. The training results of NAG vs. Enhanced NAG in the clear domain.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 2. The training results of Adagrad vs. Enhanced Adagrad in the clear domain.
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3. The training results of Adam vs. Enhanced Adam in the clear domain.
3.3.1. Experiments
Setup and Datasets We evaluate the performance of our proposed algorithms in a non-encrypted context using Python on a workstation equipped with an Intel Core G640 CPU (1.60 GHz) and 7.3 GB of RAM. As our primary objective is to accelerate training convergence, we utilize the log-likelihood function $l(\boldsymbol{\beta})$ under Maximum Likelihood Estimation (MLE) as the primary performance metric.
A comprehensive evaluation is conducted on six optimization algorithms: NAG, AdaGrad, Adam, and their respective quadratic-gradient enhanced variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam). We test these on the benchmark datasets utilized by Kim et al. (IDASH2018Andrey), which include the iDASH genomic dataset, the Myocardial Infarction dataset (edin), Low Birth Weight Study (lbw), NHANES III (nhanes3), Prostate Cancer Study (pcs), and the Umaru Impact Study (uis). The iDASH dataset, sourced from the 2017 iDASH competition (Task 3), comprises 1,579 records with 103 binary genotypes and a binary phenotype indicating cancer status.
To assess scalability, we further evaluate these algorithms on two large-scale datasets from (han2018efficient): a real-world financial dataset containing 422,108 samples and 200 features, and a restructured MNIST dataset featuring 11,982 training samples with 196 features. Throughout our experiments, we consistently employ the fixed Hessian approximation $\bar{\mathbf{H}}=-\frac{1}{4}\mathbf{X}^{\top}\mathbf{X}$ to construct the diagonal surrogate $\tilde{\mathbf{B}}$.
Hyperparameter Configuration To ensure a fair comparison with the baseline NAG (IDASH2018Andrey), Enhanced NAG adopts a learning rate of $1+\frac{10}{1+t}$, mirroring the decay structure of $\frac{10}{1+t}$ used in the original work. For Enhanced AdaGrad, we employ a modified learning rate of $N_{t}=0.06+0.01$, compared to the default $0.01$ used in standard AdaGrad. Similarly, Enhanced Adam adopts an adjusted configuration with a new augmented learning rate of $\alpha=0.011$ (calculated as the sum of a base rate $0.01$ and the original $0.001$), while maintaining the standard momentum parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. In contrast, the baseline Adam algorithm follows its default setting of $\alpha=0.001$ (kingma2014adam).
Performance Discussion As demonstrated in Figures 13, the proposed enhanced algorithms demonstrate both guaranteed convergence and a superior convergence rate compared to their first-order baselines.
Notably, the learning rates for Enhanced NAG were initially selected to maintain parity with baseline methods rather than for peak optimization. However, the framework demonstrates even more pronounced gains under alternative schedules. For instance, by adopting an exponentially decaying policy ($N_{t}=1+6\times 0.9^{t}$) on the restructured MNIST dataset, Enhanced NAG exhibits a moer clearly accelerated convergence trajectory relative to standard NAG.
Overall, the experimental results confirm that the integration of the enhanced gradient framework leads to a systematic improvement in optimization efficiency:
  • Rapid Convergence: The enhanced variants reach convergence thresholds significantly faster than their original counterparts—an essential attribute for computationally expensive environments like Homomorphic Encryption.
  • Consistent Superiority: Performance gains are consistent across diverse datasets and high-dimensional architectures, confirming the generalizability of the quadratic gradient.
  • Iteration Efficiency: Near-optimal results are typically achieved within 4 to 5 iterations, providing a critical advantage in resource-constrained or privacy-preserving settings.
Empirically, for the standalone vanilla quadratic-gradient method, optimal performance is observed when the learning rate is situated within the interval $[1,2]$. Beyond a value of $3$, the algorithm typically exhibits divergent behavior. For the Enhanced NAG variant, we recommend an exponentially decaying learning rate to synchronize rapid early-stage progress with late-stage numerical stability.
(a) First-order gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
(b) Quadratic gradient ascent: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
Figure 4. Training results of standard first-order and quadratic gradient ascent algorithms on the lbw dataset in the plaintext domain.
Gradient And Quadratic Gradient We executed the first-order vanilla gradient ascent and the proposed second-order vanilla quadratic gradient algorithms on the lbw dataset across a spectrum of learning rates, as detailed in Figure 4. These empirical results provided the foundational inspiration for the conceptualization of the quadratic gradient framework.
Crucially, our observations reveal that the quadratic gradient inherits the inherent stability and smooth progression characteristics of first-order gradients. Specifically, as the learning rate is incrementally adjusted, both the first-order gradient and second-order quadratic-gradient algorithms exhibit corresponding, continuous shifts in convergence behavior rather than erratic or abrupt fluctuations in performance.
This predictable response to learning rate modulation suggests that the quadratic gradient maintains a well-behaved optimization landscape, effectively bridging the gap between the rapid convergence of second-order methods and the robust, gradual progression of first-order optimization.
4. Secure Training
While Adaptive Gradient (AdaGrad) and Adam offer superior convergence in plaintext, they are often impractical for Homomorphic Encryption due to the high computational cost of frequent inversion and square root operations. Consequently, the Enhanced NAG method emerges as the most viable candidate for secure LR training. We leverage this method to implement an efficient, privacy-preserving LR training framework.
A key technical challenge in applying the quadratic gradient is the inversion of the diagonal matrix $\tilde{\mathbf{B}}$ to obtain $\bar{\mathbf{B}}$. To optimize performance, we offload the computation of $\bar{\mathbf{B}}$ to the data owner. Since the data owner prepares and normalizes the dataset, pre-calculating $\bar{\mathbf{B}}$ is both computationally feasible and security-neutral, as it does not leak sensitive information beyond what is already implied by the data normalization process.
4.1. Polynomial Approximation
A fundamental constraint in HE-based LR training is the inability of current homomorphic schemes to directly evaluate the non-linear Sigmoid function. To circumvent this, we approximate the Sigmoid function using a least-squares polynomial. Specifically, we utilize the 5th-degree polynomial $g(x)$ developed by Kim et al. (IDASH2018Andrey), which fits the Sigmoid function over the interval [8,8][-8,8]:
$g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}.$
4.2. Data Encoding
Given a training dataset $\mathbf{X}\in\mathbb{R}^{n\times(1+d)}$ and labels $\mathbf{Y}\in\mathbb{R}^{n\times 1}$, we follow the encoding strategy from (IDASH2018Andrey) to pack the training features and label information into a single ciphertext $\text{ct}_{Z}$. The initial weight vector $\boldsymbol{\beta}^{(0)}$ (initialized as zeros) and the diagonal entries of $\bar{\mathbf{B}}$ are replicated $n$ times to form matrices suitable for SIMD operations. The data owner then encrypts these into ciphertexts $\text{ct}_{\boldsymbol{\beta}}^{(0)}$ and $\text{ct}_{\bar{\mathbf{B}}}$, formulated as:
$\mathbf{X}$ $=\begin{bmatrix}1&x_{11}&\dots&x_{1d}\\ 1&x_{21}&\dots&x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ 1&x_{n1}&\dots&x_{nd}\\ \end{bmatrix},\quad\mathbf{Y}=\begin{bmatrix}y_{1}\\ y_{2}\\ \vdots\\ y_{n}\end{bmatrix}$,
$\text{ct}_{Z}$ $=\text{Enc}\begin{bmatrix}y_{1}&y_{1}x_{11}&\dots&y_{1}x_{1d}\\ y_{2}&y_{2}x_{21}&\dots&y_{2}x_{2d}\\ \vdots&\vdots&\ddots&\vdots\\ y_{n}&y_{n}x_{n1}&\dots&y_{n}x_{nd}\\ \end{bmatrix}$,
$\text{ct}_{\boldsymbol{\beta}}^{(0)}$ $=\text{Enc}\begin{bmatrix}\beta_{0}^{(0)}&\beta_{1}^{(0)}&\dots&\beta_{d}^{(0)}\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\dots&\beta_{d}^{(0)}\\ \vdots&\vdots&\ddots&\vdots\\ \beta_{0}^{(0)}&\beta_{1}^{(0)}&\dots&\beta_{d}^{(0)}\\ \end{bmatrix}$,
$\text{ct}_{\bar{\mathbf{B}}}$ $=\text{Enc}\begin{bmatrix}\bar{B}_{00}&\bar{B}_{11}&\dots&\bar{B}_{dd}\\ \bar{B}_{00}&\bar{B}_{11}&\dots&\bar{B}_{dd}\\ \vdots&\vdots&\ddots&\vdots\\ \bar{B}_{00}&\bar{B}_{11}&\dots&\bar{B}_{dd}\\ \end{bmatrix}$,
where $\bar{B}_{ii}$ represents the diagonal entry of $\bar{\mathbf{B}}$ derived from the fixed Hessian $\frac{1}{4}\mathbf{X}^{\top}\mathbf{X}$.
The cloud server performs the Enhanced NAG iterations on these ciphertexts to refine the encrypted weight vector. For the detailed homomorphic gradient calculation logic, we refer the reader to the methodologies in (IDASH2018Andrey).
Limitations and Trade-offs
In a privacy-preserving context, the Enhanced NAG method introduces two primary overheads compared to standard NAG: (1) one additional ciphertext-ciphertext multiplication per iteration to compute the quadratic gradient, and (2) the initial transmission of $\text{ct}_{\bar{\mathbf{B}}}$ by the data owner. However, these costs are effectively compensated by the significantly reduced number of total iterations required to reach convergence, leading to a net reduction in the overall homomorphic execution time.
Table 1. Implementation Results for iDASH datasets with 10-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
iDASH 1579 18 Ours 5 4 0.08 4.43 61.46 0.696
(IDASH2018Andrey) 5 7 0.04 6.07 62.87 0.689
Table 2. Implementation Results for other datasets with 5-fold CV
Dataset
Sample
Num
Feature
Num
Method deg $g$
Iter
Num
Storage
Learn
Time
Accuracy
AUC
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
(IDASH2018Andrey) 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
(IDASH2018Andrey) 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
(IDASH2018Andrey) 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
(IDASH2018Andrey) 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
(IDASH2018Andrey) 5 7 0.02 3.5 74.44 0.603
5. Experiments
We implement the Enhanced NAG algorithm over homomorphically encrypted data using the HEAAN library. For transparency and reproducibility, the C++ source code is publicly available at https://anonymous.4open.science/r/IDASH2017-245B. All homomorphic experiments were conducted on a public cloud instance equipped with 32 vCPUs and 64 GB of RAM.
To ensure a fair comparison with the baseline established by Kim et al. (IDASH2018Andrey), we adopt an identical 10-fold cross-validation (CV) procedure on the iDASH dataset (1,579 samples, 18 features) and 5-fold CV for the remaining five datasets. We report the average accuracy and Area Under the Curve (AUC) as the primary evaluation metrics. The experimental results, including average runtime and storage consumption (comprising the encrypted dataset for the baseline and the dataset plus $\text{ct}_{\bar{\mathbf{B}}}$ for our method), are summarized in Tables 1 and 2.
To maintain consistency in ciphertext storage, we utilize the same packing strategy proposed in (IDASH2018Andrey). This results in comparable ciphertext sizes, with the only additional overhead being the ciphertexts required to encode the diagonal approximation matrix $\bar{\mathbf{B}}$. Throughout the training procedures, we employ a dynamic learning rate of $N_{t}=1+0.9^{t}$.
The HEAAN parameters are configured to align closely with the security levels in (IDASH2018Andrey): $\log N=16$, $\log Q=1200$, $\log p=30$, and the number of slots set to $32,768$, providing a security level of $\lambda\approx 80$. Notably, we utilize a larger scaling factor, $\log p=40$, to encrypt the matrix $\bar{\mathbf{B}}$ to preserve the precision of the second-order information. Alternatively, as $\bar{\mathbf{B}}$ is derived from the data owner’s local dataset, it could theoretically be maintained in plaintext form during the cloud computation phase. By treating $\bar{\mathbf{B}}$ as a public or semi-private plaintext vector, the cloud server could utilize ciphertext-plaintext multiplication instead of the more costly ciphertext-ciphertext multiplication. This optimization would not only further reduce the computational latency per iteration but also mitigate the rate of modulus consumption, potentially allowing for additional iterations within the same multiplicative depth.
Due to the increased multiplicative depth—specifically, one additional ciphertext multiplication per iteration compared to the baseline—our scheme consumes the modulus more rapidly. Consequently, without bootstrapping, the evaluation is limited to four iterations of the Enhanced NAG algorithm. Nevertheless, as shown in the results, the proposed method achieves accuracy and AUC levels comparable to the baseline even within this restricted iteration budget.
Our learning time measurement accounts for these 4 iterations, while the baseline’s reported time includes encryption and evaluation. Given that our implementation is built upon their open-source framework, the per-iteration computational cost remains theoretically consistent with the baseline, aside from the marginal overhead of the quadratic gradient multiplication.
6. Conclusion
In this paper, we proposed a novel gradient variant, termed the quadratic gradient, and developed an enhanced NAG framework specifically tailored for training logistic regression models within the encrypted domain. By directly incorporating second-order curvature information from the Hessian matrix, the quadratic gradient effectively bridges the gap between first-order gradient-based methods and second-order Newton-type techniques.
Our approach demonstrates significant potential for accelerating a broad spectrum of optimization algorithms. Specifically, by providing a robust alternative to traditional line-search procedures, it introduces a transformative methodology for accelerating Newton-type methods through embedded gradient-descent steps. This integration maintains the structural simplicity required for homomorphic encryption while achieving superior convergence efficiency.
Future research will focus on establishing a rigorous theoretical framework to analyze the convergence guarantees of the enhanced algorithms under various conditions. Additionally, we aim to conduct extensive application-oriented evaluations across diverse learning rate configurations and high-dimensional datasets to further validate the scalability and robustness of the quadratic gradient framework.
7. Acknowledgments
The authors utilized AI-based assistance, including Grammarly (grammarly), OpenAI ChatGPT (chatgpt), and Google Gemini (gemini), to rectify typographical errors, improve grammatical structure, and refine awkward phrasing throughout the manuscript.